OpenStudy (anonymous):

1. On a certain day last year in Phoenix, AZ, it was twice as warm plus 18 degrees as it was Fargo, ND. In this linear equation, "X'' Stands for the temperature in Fargo. 2x + 18 = 70 The equation states that the temperature in Phoenix was 70 degrees. Use the equation to confirm that it was 26 degrees in Fargo. need help anyone???

OpenStudy (anonymous):

just subtract 18 from boht sides then divide by 2.

OpenStudy (anonymous):

2x=70-18 x=(70-18)/2 = 26

OpenStudy (anonymous):

your pic is really rustling my jimmies